# AI-based Multimodal Biometrics for Detecting Smartphone Distractions: Application to Online Learning

### v1

@@ 5.1 Performance for Phone Detection Models
Table 2 displays the results for the different models developed. Accuracy was selected as the primary evaluation metric due to the balanced nature of the dataset and the binary classification objective of detecting mobile phone usage. The findings suggest that using a single physiological signal leads to poorer model performance. In particular, models based on attention, meditation, alpha waves, theta waves, and delta waves exhibit the lowest accuracy. Conversely, models incorporating heart rate, beta waves, and gamma waves perform slightly better, achieving around 70% accuracy. These results indicate that mobile phone usage has a stronger impact on certain physiological signals than others.
However, individual signals were not highly effective in detecting phone usage. By combining all EEG signals with heart rate data (EEG+HR), model performance improved significantly by 8.57% (McNemar’s test, $\chi^{2}=24.45,p<0.001$), reaching an accuracy of 76%. This highlights that integrating multiple physiological signals offers a more comprehensive and reliable approach to detecting phone usage.
@@
The head pose model (HP) achieved an accuracy of 87%, improving upon the best unimodal models by 24.29% (McNemar’s test, $\chi^{2}=31.58,p<0.001$) and outperforming the model combining all EEG signals and heart rate by 14.47% (McNemar’s test, $\chi^{2}=21.02,p<0.001$). This result highlights the significance of head pose data, which often reveal notable changes in posture when interacting with a phone, compared to typical online learning behavior on a computer, making it particularly effective for detecting phone usage [9].
Nevertheless, the most effective model was the one that combines all signals (EEG+HR+HP), achieving an accuracy of 91%. This model improved upon the head pose model by 4.60% (McNemar’s test, $\chi^{2}=7.61,p=0.0058$), further confirming that integrating signals from multiple sources leads to better predictive outcomes.
It is worth noting that most models achieve the best results by smoothing the signals captured from the biosensors. This smoothing helps to reduce the noise in the data and utilizes information from a larger set of data, making the global features more relevant [17].
Figure 2 illustrates the probability density distributions of the probability scores generated by the three best models: one combining all EEG signals, heart rate, and head pose (EEG+HR+HP); another using only head pose data (HP); and the third combining all EEG signals with heart rate (EEG+HR). To assess the separation between the two classes—Phone Usage and Not Phone Usage—we calculated the Kullback–Leibler (KL) divergence. The results showed a KL divergence of $0.22$ for the EEG+HR+HP model, a divergence of $0.24$ for the HP model, and a divergence of $0.04$ for the EEG+HR model. These results indicated that the head pose-only model achieved a promising class separation performance, even when compared to a multimodal model.
However, the fact that the distributions are somewhat more separated does not necessarily mean that the performance is significantly better or worse (in fact, in terms of accuracy, combining all multimodal data achieves higher performance). In Figure 3, ROC curves are displayed for the three models, showing that depending on the threshold, either combining all multimodal data or using only head pose may be the better option. In terms of Area Under the Curve (AUC), combining all multimodal data is slightly better than using only head pose, but both models performed well overall.
Figure 3: Receiver Operating Characteristic curve (ROC) obtained for the most accurate models indicated in Table 2
@@ 5.2 Practical Implications and Limitations
While the multimodal model integrating EEG signals, heart rate, and head pose achieved the highest accuracy (91%), it is important to highlight that the head pose-based model also demonstrated strong performance. This suggests that movement patterns alone provide meaningful insights into phone usage detection.
A key advantage of the head pose model is its practicality for real-world applications. Unlike EEG or heart rate monitoring, which require specialized and intrusive biosensors, head pose estimation can be performed using standard webcams, making it a more accessible and scalable solution. This opens the door for integration into online learning platforms that already track head movements or engagement levels. By incorporating head pose analysis, these platforms could provide real-time feedback to learners, helping them stay focused and minimizing phone-induced distractions.
Given the application limitations of the model that combines all biometric signals, future work could explore estimating these variables through the webcam. If these estimations prove reliable, it would be possible to integrate a proxy version of the multimodal model without requiring additional wearable sensors. This approach would not only maintain ease of implementation and user comfort but also enhance the generalizability of the models by enabling data collection through more accessible and cost-effective methods. By eliminating the need for expensive biosensors, this solution could facilitate the widespread adoption of distraction detection systems in online learning environments, making advanced analytics available to a broader range of users.
Although the number of samples in the IMPROVE dataset is not exceptionally large (132 instances), it is still considered a relatively substantial dataset within the MMLA field. This study has demonstrated that it is indeed feasible to develop accurate models for detecting smartphone usage using multimodal data. As a result, this work provides an empirical basis for future research efforts aiming to scale and refine these models. Subsequent studies may focus on increasing the number of participants to improve model generalizability, potentially by relying solely on more accessible modalities such as head pose estimation.
@@ 6 Conclusion and Future Work
This article explores the potential of MMLA to enhance the detection of phone usage among learners by leveraging various biometric variables. Several AI-based models were proposed and evaluated to predict when learners use their phones and become distracted during an online course. Specifically, both unimodal models (using individual biometric variables) and multimodal models (combining multiple variables) were developed.
After comparing several algorithms, the results indicate that unimodal models relying on a single EEG signal or heart rate performed significantly worse than the model that integrated all EEG signals and heart rate. The second-best model utilized head pose data, suggesting that movement patterns provide valuable information about phone usage. However, the highest-performing model incorporated all multimodal data (EEG signals, heart rate, and head pose), achieving an accuracy of 91%. This supports the idea that using multiple biosensors to capture different biometric signals leads to more accurate prediction models. These findings align with previous MMLA research, which highlights the value of multimodality in improving the performance of prediction models across various learning tasks [22, 37].
As future work, we will focus on other types of distractions, such as browsing or checking emails, and we will aim to develop models that can detect a wider variety of learning events. Additionally, we will investigate other indicators that have proven valuable for detecting learner behaviors, such as eyeblink [18], pupil dilation [27], eye gaze patterns [32], and typing behavior [1], among others. Furthermore, we will examine the integration of both local and global features [17] and we will attempt to extend this research to face-to-face learning with multiple students in order to develop more general models.
Finally, we plan to explore the use of real-time interventions that notify learners when they are becoming distracted, helping them stay focused and to evaluate the impact of such interventions on students’ learning outcomes. Additionally, we will evaluate the effectiveness of these interventions by examining their impact on learners’ learning outcomes, including engagement, comprehension, and overall performance. Through this research, we hope to gain a deeper understanding of how personalized feedback can enhance the learning experience in online courses.
Furthermore, the methods developed here are useful in estimating distractions that arise from the undesirable use of smartphones in any monitorable task that requires concentration, not only computer-based online learning. In our future work we will also explore how our methods apply to areas like autonomous driving (in which the driver must pay full attention to the road sometimes [24]) or traditional book-and-paper study).

### v2

@@ 5.1 Performance for Phone Detection Models
Table 2 displays the results for the different models developed. Accuracy was selected as the primary evaluation metric due to the balanced nature of the dataset and the objective of binary classification to detect mobile phone usage. The findings suggest that the use of a single physiological signal leads to poorer model performance. In particular, models based on attention, meditation, alpha waves, theta waves, and delta waves exhibit the lowest accuracy. In contrast, models that incorporate heart rate, beta waves, and gamma waves perform slightly better, achieving around 70% accuracy. These results indicate that mobile phone usage has a stronger impact on certain physiological signals than others.
However, individual signals were not highly effective in detecting phone usage. By combining all EEG signals with heart rate data (EEG + HR), the model performance improved significantly by 8.57% (McNemar’s test, $\chi^{2}=24.45,p<0.001$), reaching an accuracy of 76%. This highlights that integrating multiple physiological signals offers a more comprehensive and reliable approach to detecting phone usage.
@@
The head pose model (HP) achieved an accuracy of 87%, improved on the best unimodal models by 24.29% (McNemar’s test, $\chi^{2}=31.58,p<0.001$) and outperformed the model combining all EEG signals and heart rate by 14.47% (McNemar’s test, $\chi^{2}=21.02,p<0.001$). This result highlights the importance of head pose data, which often reveal notable changes in posture when interacting with a phone, compared to typical online learning behavior on a computer, making it particularly effective in detecting phone usage [11].
However, the most effective model was the one that combined all signals (EEG + HR + HP), achieving an accuracy of 91%. This model improved on the head pose model by 4.60% (McNemar’s test, $\chi^{2}=7.61,p=0.0058$), further confirming that integrating signals from multiple sources leads to better predictive outcomes.
It should be noted that most models achieve the best results by smoothing the signals captured from the biosensors. This smoothing helps reduce the noise in the data and utilizes information from a larger set of data, making global features more relevant [19].
Figure 2 illustrates the probability density distributions of the probability scores generated by the three best models: one that combines all EEG signals, heart rate, and head pose (EEG+HR+HP); another using only head pose data (HP); and the third combining all EEG signals with heart rate (EEG+HR). To assess the separation between the two classes, Phone usage and Not Phone usage, we calculated the Kullback-Leibler (KL) divergence. The results showed a KL divergence of $0.22$ for the EEG+HR+HP model, a divergence of $0.24$ for the HP model and a divergence of $0.04$ for the EEG+HR model. These results indicated that the head pose-only model achieved promising class separation performance, even when compared to a multimodal model.
However, the fact that the distributions are somewhat more separated does not necessarily mean that the performance is significantly better or worse (in fact, in terms of accuracy, combining all multimodal data achieves higher performance). In Figure 3, the ROC curves are displayed for the three models, showing that depending on the threshold, combining all multimodal data or using only the head pose may be the best option. In terms of Area Under the Curve (AUC), combining all multimodal data is slightly better than using only the head pose, but both models performed well overall.
Figure 3: ROC obtained for the most accurate models indicated in Table 2
@@ 5.2 Practical Implications and Limitations
Although the multimodal model integrating EEG signals, heart rate, and head pose achieved the highest accuracy (91%), it is important to note that the head pose-based model also demonstrated strong performance. This suggests that movement patterns alone provide meaningful insight into phone usage detection.
A key advantage of the head pose model is its practicality for real-world applications. Unlike EEG or heart rate monitoring, which may require specialized and intrusive biosensors [30], estimation of head posture can be performed using standard webcams, making it a more accessible and scalable solution. This opens the door to integration into on-line learning platforms that already track head movements or engagement levels. By incorporating head pose analysis, these platforms could provide real-time feedback to learners, helping them stay focused and minimizing phone-induced distractions.
Given the application limitations of the model that combines all biometric signals, future work could explore estimating these variables through the webcam. If these estimations are reliable, it would be possible to integrate a proxy version of the multimodal model without requiring additional wearable sensors. This approach would not only maintain ease of implementation and user comfort, but would also enhance the generalizability of the models by enabling data collection through more accessible and cost-effective methods. By eliminating the need for expensive biosensors, this solution could facilitate the widespread adoption of distraction detection systems in online learning environments, making advanced analytics available to a broader range of users.
Although the number of samples in the IMPROVE dataset is not exceptionally large (132 instances), it is still considered a relatively substantial dataset within the MMLA field. This study has shown that it is feasible to develop accurate models to detect smartphone usage using multimodal data. As a result, this work provides an empirical basis for future research efforts to scale and refine these models. Subsequent studies may focus on increasing the number of participants to improve the generalizability of the model, potentially relying solely on more accessible modalities such as the estimation of the head pose.
@@ 6 Conclusion and Future Work
This article explores the potential of MMLA to enhance the detection of phone usage among learners by using various biometric variables. Several AI-based models were proposed and evaluated to predict when learners use their phones and become distracted during an online course. Specifically, both unimodal models (using individual biometric variables) and multimodal models (combining multiple variables) were developed.
After comparing several algorithms, the results indicate that unimodal models relying on a single EEG signal or heart rate performed significantly worse than the model that integrated all EEG signals and heart rate. The second-best model used head pose data, suggesting that movement patterns provide valuable information about phone usage. However, the highest-performing model incorporated all multimodal data (EEG signals, heart rate, and head pose), achieving an accuracy of 91%. This supports the idea that using multiple biosensors to capture different biometric signals leads to more accurate prediction models. These findings align with previous MMLA research, which highlights the value of multimodality [25, 41] in improving the performance of prediction models on various learning tasks [28, 45].
In future work, we will focus on other distractions, such as browsing or checking emails, and we will aim to develop models that can detect a wider variety of learning events. In addition, we will investigate other indicators that have proven valuable for detecting learner behaviors, such as eyeblink [20], pupil dilation [34], eye gaze patterns [39], typing [2, 46], mouse [1] and touchscreen dynamics [26, 23], among others. Furthermore, we will examine the integration of both local and global features [19] and will attempt to extend this research to face-to-face learning with multiple students in order to develop more general models.
Finally, we plan to explore real-time interventions that notify learners when they are becoming distracted, helping them stay focused, and to evaluate the impact of such interventions on students’ learning outcomes. In addition, we will evaluate the effectiveness of these interventions by examining their impact on learners’ learning outcomes, including engagement, comprehension, and overall performance. Through this research, we hope to gain a deeper understanding of how personalized feedback can enhance the learning experience in online courses.
Furthermore, the methods developed here are useful in estimating distractions that arise from the undesirable use of smartphones in any monitorable task that requires concentration, not only computer-based online learning. In our future work, we will also explore how our methods apply to areas such as autonomous driving (in which the driver must pay full attention to the road sometimes [31]) or traditional book-and-paper study.